OpenStudy (anonymous):

et A and B be 2*2 matrices such that 2A+3B= 1 2 -1 -2 Atranspose +B transpose=-1 3 -7 6 Compute A and B

OpenStudy (anonymous):

one sec I'll do it real quick

OpenStudy (anonymous):

maybe not quickly

OpenStudy (anonymous):

:D take ur time:)

OpenStudy (anonymous):

ooooo right bad at addition, woops

OpenStudy (anonymous):

u can attach a picture of ur paper with solution

OpenStudy (anonymous):

check to see if these work correctly: A = |-4 -23| |16 -5|

OpenStudy (anonymous):

B = |3 26| |-7 8|

OpenStudy (anonymous):

may u send me ur steps also

OpenStudy (anonymous):

nono B = |3 16| |-7 8|

OpenStudy (anonymous):

umm it's fairly illegible

OpenStudy (anonymous):

just attach a picture with ur steps and btw its perfectly right:)

OpenStudy (anonymous):

the camera on my phone is pretty terrible, can't read my writing in pencil

OpenStudy (anonymous):

steps I used: A^T + B^T = (A+B)^T thus: a11 + b11 = -1 a22+ b22 = 6 a12 + b12 = -7 a21 + b21 = 3 and i solved for the a values and plugged them into the other set of equations

OpenStudy (anonymous):

which are: 2a11 + 2b11 = 1 2a12 + 3b12 = 2 2a21 + 3b21 = -1 2a22+ 3b22 = -2

OpenStudy (anonymous):

which gave me the values for B matrix then I plugged those values back into the first set of equations to get my A matrix

OpenStudy (anonymous):

thanks alot:)

OpenStudy (anonymous):

yep, matricies are fun!

OpenStudy (anonymous):

wants alot of them?:D

OpenStudy (anonymous):

hit me with another one, I'm tired of my linear algebra studying anyway

OpenStudy (anonymous):

Let A be a 2*2 matrix Prove that if det(a)not equal zero then the inverse is given by A^-1=(1/det(A))adj(A)

OpenStudy (anonymous):

check that alsoo s

OpenStudy (anonymous):

hmm, i konw it has to do with the cofactor expansion, let me chew on that a sec

OpenStudy (anonymous):

err, cofactor matrix

OpenStudy (anonymous):

??

OpenStudy (anonymous):

thinking on it, I should really know this.

OpenStudy (anonymous):

ok :) take ur time

OpenStudy (anonymous):

from cramer's rule we get:\[x_{j} = \det(A _{i}(e _{j})) / \det(A)\] thus \[x _{ij} = (1/\det(A))C _{ji}\] which gives us:\[A^{-1} = X = (1/\det(A))[C _{ji}] = (1/\det(A))[C _{ij}]^T\]

OpenStudy (anonymous):

and \[[C _{ij}]^T = adj(A)\]

OpenStudy (anonymous):

might need an Assume the \[\det(A) \ne 0\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i pieced together a few statements, I'm pretty sure that is the right aproach but can't tell you how to do it exactly

OpenStudy (anonymous):

and about the attachment did u check it? may u please check it:)

OpenStudy (anonymous):

couldn't load it

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cramer's_rule

OpenStudy (anonymous):

thanks:) i will send u the attachment again

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yeah sorry lost the connection

OpenStudy (anonymous):

thats the Gauss Jordan algorithm metod whats about the adjugates

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

wouldn't let me type, look at both

OpenStudy (anonymous):

perfect:)

OpenStudy (anonymous):

want more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

start a new thread though

OpenStudy (anonymous):

ohh sorry ok:)